Zorba Zantani runs a business, Zantani Manufacturers. Zantani Manufacturers, manufactures kitchen utensils and sells them to retailers who, in turn, sells the product to the public.

Zantani Manufacturers is a registered VAT vendor and only deals with other VAT vendors.

All amounts include VAT at 15%, where applicable, unless otherwise indicated.

In February 2019, Zantani Manufacturers manufactured 1 500 potato peelers at a cost of R690 each (including VAT).

The labour cost to produce the 1 500 potato peelers amounted to R25 000 .








4. Calculate the total sales amount (including VAT) of the 1 500 potato peelers. (5)

5. Calculate the cost of sales incurred by Outside Retailers. (1)

6. Calculate Zantani Manufacturers gross margin. (Round to the nearest whole percentage) (3)



1
Expert's answer
2021-05-13T17:56:05-0400

Solution:

4.). Total sales amount inclusive of VAT = Selling Price x Quantity sold

Quantity sold = 1,500 potato peelers

Selling price = Cost of production + Mark-up

130% = 100% + 30%

Therefore, selling price = 130% of the cost = 130% "\times" 690 = R897 (VAT inclusive)

Total sales = 897 x 1,500 = R1,345,500

Total sales amount (inclusive of VAT) = R1,345,500

 

5.). The cost of sales incurred by Outside Retailers = Mark-up charged

The cost of sales = Total sales – Total Cost of production

The cost of sales = (897x1500) – ((690x1500) + 25,000

The cost of sales = 1,345,500 – 1,060,000 = 285,500

The cost of sales incurred by outside retailers = R285,500

 

6.). Zantani manufacturer’s gross margin = "\frac{Sales \; revenue - cost\; of\; sales}{sales\; revenue\; }\times 100"


Gross margin = "\frac{1,345,500 - 1,060,000}{1,345,500}\times 100 = 21.22\%"

Gross margin = 21.22%